OpenStudy (anonymous):

guys ,,, can u answer this question double differentiate xy-y^2=10

OpenStudy (amistre64):

its not so much an answer as it is a process

OpenStudy (amistre64):

and i assume you mean you want to derive it twice?

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

do you know how to derive? i mean the rules such as power rule, product rule, chain rule etc?

OpenStudy (anonymous):

i think i should use product rule

myininaya (myininaya):

why do u always say derive amistre lol

OpenStudy (amistre64):

this takes both a product rule and a power rule in the first derivative

OpenStudy (amistre64):

:) its how i make sense of the sale on Dwords that calculus bought

myininaya (myininaya):

lol

OpenStudy (amistre64):

d (xy-y^2=10) --- dx d (xy) d (-y^2) d (10) --- + ---- = ----- dx dx dx

OpenStudy (anonymous):

first i got -y/(x+2y)

myininaya (myininaya):

(xy)'=xy'+y (-y^2)'=-2yy'

myininaya (myininaya):

(constant)'=0

OpenStudy (amistre64):

dx dy dx --- (y) + --- (-2y) = -- (0) ; and dx/dx = 1 dx dx dx

OpenStudy (amistre64):

.... yeah, i messed up that xy, forgot to do a product rule :)

OpenStudy (amistre64):

dx dy dy dx --- (y) + --- (x) + --- (-2y) = -- (0) ; and dx/dx = 1 dx dx dx dx thats better

OpenStudy (amistre64):

dy dy y + --- (x) + --- (-2y) = 0 dx dx

OpenStudy (anonymous):

it will be dy/dx=-y/x+2y

OpenStudy (anonymous):

right

OpenStudy (amistre64):

derive it again i spose d (y + y'x + y' -2y = 0) --- dx d (y) d (xy') d (-2y y') d (0) --- + --- + --- = --- dx dx dx dx dy dx d2y d2y dy dx --- (1) + --- (y') +----(x) + ---- (-2y) + --- (-2y') = --- (0) dx dx dx2 dx2 dx dx

myininaya (myininaya):

i would had solved for y' in the first one but lets see what happens with amistre;s approach

OpenStudy (amistre64):

y' + y' + y'' x +y'' -2y + y' -2y' = 0 2y'+ y''(x -2y) -2(y'^2) = 0 y''(x -2y) = 2(y'^2) - 2y' y'' = 2y'(y'-1)/(x-2y)

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

since: y' = -y/(x-2y) we get... \[y'' = \frac{-2y}{(x-2y)}*\frac{\frac{-y}{(x-2y)} -1}{(x-2y)}\] \[y'' = \frac{-2y}{(x-2y)}*\frac{\frac{-y-(x-2y)}{(x-2y)}}{(x-2y)}\] \[y'' = \frac{-2y}{(x-2y)}*\frac{-y-x+2y}{(x-2y)^2}\] \[y'' = \frac{-2y}{(x-2y)}*\frac{y-x}{(x-2y)^2}\] \[y'' = \frac{-2y^2+2xy}{(x-2y)^3}\] perhaps?

OpenStudy (amistre64):

prolly messed it up along the way :)

myininaya (myininaya):

i think it looks good

OpenStudy (amistre64):

then I do too :)

myininaya (myininaya):

:)

myininaya (myininaya):

wait i dont know

myininaya (myininaya):

\[xy-y^2=10\] \[xy'+y-2yy'=0\] \[y'(x-2y)=-y\] \[y'=\frac{-y}{x-2y}\] \[y''=\frac{-y'(x-2y)-(1-2y')(-y)}{(x-2y)^2}\] \[y''=\frac{\frac{y}{x-2y}(x-2y)-(1+\frac{2y}{x-2y})(-y)}{(x-2y)^2}\] \[y''=\frac{y+y+\frac{2y^2}{x-2y}}{(x-2y)^2}\] \[y''=\frac{2y(x-2y)+2y^2}{(x-2y)^3}\] \[y''=\frac{2xy-2y^2+2y^2}{(x-2y)^3}\] \[y''=\frac{2xy}{(x-2y)^3}\] maybe i made a mistake though

myininaya (myininaya):

oops i did you did good amistre

myininaya (myininaya):

see 2y(x-2y)=2xy-4y^2 and i put 2xy-2y^2

myininaya (myininaya):

\[y''=\frac{2xy-4y^2+2y^2}{(x-2y)^3}=\frac{2xy-2y^2}{(x-2y)^3}\]

OpenStudy (amistre64):

yeah, i missed it on the algebra lol

myininaya (myininaya):

no you did good and i did bad or do you mean you didn't see my algebraic mistake?

OpenStudy (amistre64):

i was going thru your work thinking it was mine :)

OpenStudy (amistre64):

but i see now we got the same results :)

OpenStudy (amistre64):

we apparently have the same handwriting

myininaya (myininaya):

lol

myininaya (myininaya):

it looks similar

myininaya (myininaya):

we do have different writing styles though you use dy/dx notation and i use y'

OpenStudy (amistre64):

i like to use d/dx to show the background movements of implicit differentation

OpenStudy (amistre64):

people seem to have a notion that it is some magical conflageration that is somehow different from what they learn derivatives to be

myininaya (myininaya):

i like y' because i don't like writing fractions it is just easier for me to write y' then to write dy/dx

OpenStudy (amistre64):

y' cleans it up nice for me

myininaya (myininaya):

either notation is fine but i think dy/dx notation is suppose to be better and i cant remember why

OpenStudy (amistre64):

pros and cons, thats all

myininaya (myininaya):

i think they said Europe or some country/continent was behind in math because they never used dy/dx notation i;m sure but im just recalling false stuff

myininaya (myininaya):

maybe im recalling false stuff*

OpenStudy (amistre64):

thats right; england stuck to newtons notation (y') while the rest of the europe used leibniz (dy/dx) and england fell behind

myininaya (myininaya):

right y' is newton's and dy/dx was leibniz you're so smart :)

myininaya (myininaya):

i would had fell behind with england then

myininaya (myininaya):

amistre=the rest of europe

OpenStudy (amistre64):

us sicilians never worried to much about it ;)

myininaya (myininaya):

i want their pizza